OpenStudy (anonymous):

solve the rational equation. check your answer. 5/ 2x - 3 = 7/3x

OpenStudy (anonymous):

15x=14x-21 do the equation

OpenStudy (anonymous):

x=-21

OpenStudy (tkhunny):

\(\dfrac{5}{2x-3} = \dfrac{7}{3x}\) Is that it? #1 Domain Issues. \(2x-3\ne 0\;and\;x\ne0 \) Now what?

OpenStudy (anonymous):

@tkhunny there's no domaine issue

OpenStudy (tkhunny):

#1 Domain Issues. #2 Anyone who says there are no Domain Issues is being careless. Now what?

OpenStudy (anonymous):

The final answer should be -21, I just need to know how to show my work.

OpenStudy (tkhunny):

You show your work by showing your work. You must not EVER multiply by ANYTHING unless you KNOW it is NOT zero. This is why there are Domain Issues. \(\dfrac{5}{2x-3} = \dfrac{7}{3x}\) In this form, it should be clear that a value x = 0 is unacceptable. Once we say that, we can multiply by x, giving. \(\dfrac{5x}{2x-3} = \dfrac{7}{3}\) 3 is never zero, so we can multiply by that. \(\dfrac{15x}{2x-3} = 7\) In this form, it should be clear that 2x-3 = 0 is not acceptable. Once we say \(x\ne 3/2\), we can multiply by 2x-3, giving \(15x = 7(2x-3)\) Unfortunately, if I keep typing, you will only be able to show MY work. You do the rest.

OpenStudy (anonymous):

I don't know how to continue

OpenStudy (jhannybean):

I think what @tkhunny was implying is that look at all parts of the problem, ifyou have a variable in the denominator find out what it means when they mentioned that \(\large x \neq \frac{ 3 }{ 2 }\) that means there will be a vertical asymptote at (3/2), hence the "domain issue" so now we're given \[15x=7(2x-3) \]\(15x=14x-21\) solve for x

OpenStudy (anonymous):

Thank you

OpenStudy (jhannybean):

No problem